Create an equation of a line in y=mx+b form that is perpendicular to the line:
y= (-1/4)x+180 and goes through the point (8,12)

Answer by ewatrrr(24785)

Line perpendicular to the line y= (-1/4)x+180
m = 4
|Using P(8,12)
y - 12 = 4(x-8)
y - 12 = 4x-32
y = 4x -20